Required information

[The following information applies to the questions displayed below.]

Green Grow Inc. (GGI) manufactures lawn fertilizer. Because of the product’s very high quality, GGI often receives special orders from agricultural research groups. For each type of fertilizer sold, each bag is carefully filled to have the precise mix of components advertised for that type of fertilizer. GGI’s operating capacity is 33,000 one-hundred-pound bags per month, and it currently is selling 31,000 bags manufactured in 31 batches of 1,000 bags each. The firm just received a request for a special order of 7,200 one-hundred-pound bags of fertilizer for $200,000 from APAC, a research organization. The production costs would be the same, but there would be no variable selling costs. Delivery and other packaging and distribution services would cause a one-time $4,400 cost for GGI. The special order would be processed in two batches of 3,600 bags each. (No incremental batch-level costs are anticipated. Most of the batch-level costs in this case are short-term fixed costs, such as salaries and depreciation.) The following information is provided about GGI’s current operations:

Sales and production cost data for 31,000 bags, per bag:
Sales price $ 46
Variable manufacturing costs 12
Variable selling costs 3
Fixed manufacturing costs 21
Fixed marketing costs 7

No marketing costs would be associated with the special order. Because the order would be used in research and consistency is critical, APAC requires that GGI fill the entire order of 7,200 bags.

Required:

1. What is the total relevant cost of filling this special sales order?

2. What would be the change in operating income if the special order is accepted?

3. What is the break-even selling price per unit for the special sales order (i.e., what is the selling price that would result in a zero effect on operating income)?

4. Prepare comparative income statements, using the contribution format, for both the current situation and assuming the special order is accepted at the break-even price determined in requirement 3.
Answer #1
Part-1
computation of Relevant cost for special Order
Relevant manufacturing and distribution cost = Variable manufacturing cost * Number of bags + One time distribution cost
= 7200 * $12 + $4400= $90800
computation of Opportunity Cost
Excess Capacity to be produced ( 7200-(33000-31000)=5200
Opportunity cost = 5200 * (Profit per bag) = 5200 * (46-12-3) = $161200
computation of Total Relevant Cost for Special order
Relevant cost for special Order:- $90800
Opportunity Cost for Special Order: $161200

Total relevant cost : $252000

2.Change in operating income = Relevant cost - Selling price of special order = $252000- 200000 = $52000

3. Break even selling price = Total relevant Cost / Special order = 252000 /7200 = $35
4. Comparitive Income Statement
Current Situation New Situation
Regular Special Total
Sales Unit 31000 25800 7200 33000
Regular Order @46 $14,26,000 $11,86,800 $11,86,800
Special Order @35 $2,52,000 $2,52,000
Less: Variable Cost
Manufaccturing @$12 $3,72,000 $3,09,600 $86,400 $3,96,000
Marketing $3 $93,000 $77,400 $77,400
Contribution Margin $9,61,000 $9,65,400
Less: Fixed Cost
Manufacturing @21 $6,51,000 $6,51,000 $6,51,000
Marketing @7 $2,17,000 $2,17,000 $2,17,000
One Time Payment $4,400 $4,400
Operating Income $93,000 $93,000

Kindly rate positively it would help me lot....